OpenStudy (anonymous):

Which set of equations would be used to solve this word problem? An airplane flew 5 hours with a 50 mph head wind. The return trip with a tail wind of the same speed took 3 hours. Find the speed of the plane in still air. d = 8(r + 50) d = 5(r + 50) and d = 3(r − 50) d = 5(r − 50) and d = 3(r + 50) None of these systems can be used to solve this problem.

OpenStudy (anonymous):

d = 5(r − 50) and d = 3(r + 50) 5 hours = rate -50mph winds thats answer..

OpenStudy (radar):

d=5(r-50) d=3(r+50) d=d 5r-250=3r+150 2r=400 r=200 mph d=750 miles